OpenStudy (anonymous):

prove the derivative of cos^-1(x) is -1 / radical(1-x^2)?

OpenStudy (jamesj):

Let y = arccos(x). Then x = cos(y) Now differente both sides ofthis last relation and use implicit differentiation. \[\frac{d \ }{dx} x = \frac{d \ }{dx} \cos y = \frac{dy}{dx} \frac{d \ }{dy} \cos y\]

OpenStudy (jamesj):

Now the left hand side, LHS = 1 and the right RHS = dy/dx . (-sin y) = dy/dx . - sqrt(1-x^2)

OpenStudy (jamesj):

Hence \[1 = \frac{dy}{dx} .( - \sqrt{1-x^2}) \implies \ \frac{dy}{dx} = \frac{-1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

you can use implicit diff, or the chain rule \[x=\cos(\arccos(x))\] \[1=-\sin(\arccos(x))\times \arccos'(x)\] \[\arccos'(x)=-\frac{1}{\sin(\arccos(x))}\] and then you have to think back to trig as to how to find \[\sin(\arccos(x))=\sqrt{1-x^2}\] although that should be easy